# Self-Learned Representation-Guided Latent Diffusion Model for Breast Cancer Classification in Deep Ultraviolet Whole Surface Images

### v1

@@ Abstract
Breast-Conserving Surgery (BCS) requires precise intraoperative margin assessment to preserve healthy tissue. Deep Ultraviolet Fluorescence Scanning Microscopy (DUV-FSM) offers rapid, high-resolution surface imaging for this purpose; however, the scarcity of annotated DUV data hinders the training of robust deep learning models. To address this, we propose an Self-Supervised Learning (SSL)-guided Latent Diffusion Model (LDM) to generate high-quality synthetic training patches. By guiding the LDM with embeddings from a fine-tuned DINO teacher, we inject rich semantic details of cellular structures into the synthetic data. We combine real and synthetic patches to fine-tune a Vision Transformer (ViT), utilizing patch prediction aggregation for WSI-level classification. Experiments using 5-fold cross-validation demonstrate that our method achieves 96.47% accuracy and reduces the FID score to 45.72, significantly outperforming class-conditioned baselines.
@@ 1 Introduction
Breast cancer affects one in eight women in the United States, making Breast-Conserving Surgery (BCS) a critical procedure for removing tumors while preserving healthy tissue [b1, b2]. Deep Ultraviolet Fluorescence Scanning Microscope (DUV-FSM) produces high-resolution, real-time images of tissue surfaces, enabling easier distinction between cancerous and healthy areas. When combined with automated classification, it can help improve margin assessment and reduce extra surgeries. Classifying Deep Ultraviolet Whole Surface Images (DUV WSIs) at full resolution is impractical and downsampling can lose important details. To address this, patch-level Deep Learning (DL) methods have been used [b1, b2, b4]. Despite their benefits, deep learning models often don’t perform well with limited data, which is common in medical imaging, such as DUV. This shows the need for good data augmentation and stronger learning methods to improve their performance [b5, b6].
While Generative Adversarial Networks (GANs) [b7] are widely used for generating synthetic data, they often produce limited diversity, and suffer from mode collapse. This is particularly problematic for unbalanced DUV WSI patch-level datasets that require fine-grained malignant samples [b5, b8, b12]. Denoising Diffusion Probabilistic Models (DDPMs) [b22] offer higher stability and image quality, especially for DUV WSI data [b5], but entail high computational costs [b13]. LDMs [b15] reduce computation by performing diffusion in a lower-dimensional latent space. Their cross-attention module enables guided image generation using different forms of conditioning information [b15]. Recent works [b16, b17] have shown that conditioning diffusion models on Self-Supervised Learning (SSL) embeddings, rather than text or simple class labels, yields semantically richer images.

### v2

@@ Abstract
Breast-Conserving Surgery (BCS) requires precise intraoperative margin assessment to preserve healthy tissue. Deep Ultraviolet Fluorescence Scanning Microscopy (DUV-FSM) offers rapid, high-resolution surface imaging for this purpose; however, the scarcity of annotated DUV data hinders the training of robust deep learning models. To address this, we propose an Self-Supervised Learning (SSL)-guided Latent Diffusion Model (LDM) to generate high-quality synthetic training patches. By guiding the LDM with embeddings from a fine-tuned DINO teacher, we inject rich semantic details of cellular structures into the synthetic data. We combine real and synthetic patches to fine-tune a Vision Transformer (ViT) and use patch-prediction aggregation for WSI-level classification. Experiments using 5-fold cross-validation demonstrate that our method achieves 96.47% accuracy and reduces the FID score to 45.72, significantly outperforming class-conditioned baselines.
@@ 1 Introduction
Breast cancer affects one in eight women in the United States, making Breast-Conserving Surgery (BCS) a critical procedure for removing tumors while preserving healthy tissue [b1, b4]. Deep Ultraviolet Fluorescence Scanning Microscope (DUV-FSM) produces high-resolution, real-time images of tissue surfaces, enabling easier distinction between cancerous and healthy areas. When combined with automated classification, it can help improve margin assessment and reduce extra surgeries. Classifying Deep Ultraviolet Whole Surface Images (DUV WSIs) at full resolution is impractical and downsampling can lose important details. To address this, patch-level Deep Learning (DL) methods have been used [b1, b4]. Despite their benefits, deep learning models often don’t perform well with limited data, which is common in medical imaging, such as DUV. This shows the need for good data augmentation and stronger learning methods to improve their performance [b5, b6].
While Generative Adversarial Networks (GANs) [b7] are widely used for generating synthetic data, they often produce limited diversity and suffer from mode collapse. This is particularly problematic for unbalanced DUV WSI patch-level datasets that require fine-grained malignant samples [b5, b8, b12]. Denoising Diffusion Probabilistic Models (DDPMs) [b22] offer higher stability and image quality, especially for DUV WSI data [b5], but entail high computational costs [b13]. LDMs [b15] reduce computation by performing diffusion in a lower-dimensional latent space. Their cross-attention module enables guided image generation using different forms of conditioning information [b15]. Recent works [b16, b17] have shown that conditioning diffusion models on Self-Supervised Learning (SSL) embeddings, rather than text or simple class labels, yields semantically richer images.